OpenStudy (anonymous):

find the gcf of monomials 27x^2 and 45x^3 y^2

OpenStudy (watchmath):

\(27x^2=3(9x^2)\) \(45x^3y^2=5xy^2(9x^2)\) So \(9x^2\) is the gcf.